# Boundary layers, transport and universal distribution in boundary driven active systems

### v2

@@ Appendix A Exit probability of RTP at the left boundary
$E_{m}(x)=\frac{(1+e^{-\mu L})\,x-\frac{v}{2\,\omega}(m-\frac{v}{\mu\,D})\,[1-e^{-\mu x}]+\frac{v}{2\,\omega}(m+\frac{v}{\mu\,D})\,[e^{-\mu(L-x)}-e^{-\mu L}]}{L\,(1+e^{-\mu L})+\frac{v^{2}}{\omega\,\mu\,D}\,(1-e^{-\mu L})}.$ (A.2)
@@
Appendix B Expression for the left eigenvectors of $\mathcal{L}$ in the $\lambda=0$ band and the distribution
We find the left eigenvectors $\langle\psi_{n}|$ of $\mathcal{L}$ by noting that it is just the transpose of the right eigenvector of $\mathcal{L}^{\dagger}\equiv\mathcal{L}|_{\partial_{x}\rightarrow-\partial_{x}}$:
$\langle\psi_{n}(x)|$ $=$ $|\phi_{n}(x,-v)\rangle^{T}\approx\alpha^{\prime}_{n}\bigg{(}2\,i\,\sin\frac{n\pi x}{L}\begin{bmatrix}1\\ 1\end{bmatrix}^{T}+\,i\,\frac{n\pi v}{\omega\,L}\cos\frac{n\pi x}{L}\begin{bmatrix}1\\ -1\end{bmatrix}^{T}$ (B.1)
@@ Appendix B Expression for the left eigenvectors of $\mathcal{L}$ in the $\lambda=0$ band and the distribution
[…]
@@
$\alpha^{\prime}_{n}\alpha_{n}\approx-\frac{1}{4L}\left(1+i(i+n\pi)\,\frac{2\,l_{M}}{L}\right).$ (B.3)
For the initial condition, $|P(x,0)\rangle=\delta(x-x_{0})\,\begin{bmatrix}s_{+}\\ s_{-}\end{bmatrix}$, we find,
$\begin{split}c_{n}&=\int_{0}^{L}dx\,\langle\psi_{n}(x)|P(x,0)\rangle\approx\alpha^{\prime}_{n}\bigg{[}2i\sin\frac{n\pi x_{0}}{L}+\frac{in\pi v}{\omega L}(s_{+}-s_{-})\cos\frac{n\pi x_{0}}{L}\\ &~{}~{}~{}~{}~{}-\frac{n\pi v}{\omega L}(1-\frac{v}{\sqrt{2\omega D_{e}}})\sin\frac{n\pi x_{0}}{L}+\frac{in\pi v^{2}}{\omega L\sqrt{2\omega D_{e}}}(1-\frac{2x_{0}}{L})\cos\frac{n\pi x_{0}}{L}\\ &~{}~{}~{}~{}~{}-\frac{in\pi v}{\omega L}(1+\frac{v}{\sqrt{2\omega D_{e}}})\bigg{\{}(-1)^{n-1}e^{-k_{a}(L-x_{0})}(\tilde{a}_{n}^{(1)}s_{+}+s_{-})+e^{-k_{a}x_{0}}(s_{+}+\tilde{a}_{n}^{(1)}s_{-})\bigg{\}}\bigg{]}.\end{split}$ (B.4)
Using Eqs. (B.3) and (B.4) in Eq. (62), we obtain Eq. (64) for the propagator to $O(L^{-1})$. To find a closed-form expression for times $\omega^{-1}\ll t\ll L^{2}/D_{e}$, we define the following quantities corresponding to the summations that occur in Eq. (64) and use the Poisson summation formula:
@@
The $L$-dependent terms in the square brackets, which are very small for large $L$, are kept in view of the symmetries satisfied by the discrete summation. These also preserve the symmetries of the density and magnetisation profile:
$\begin{split}P(x,t;x_{0},m_{0})&=P(L-x,t;L-x_{0},-m_{0}),\\ Q(x,t;x_{0},m_{0})&=-Q(L-x,t;L-x_{0},-m_{0}).\end{split}$ (B.6)
Using Eq. (B.5) we rewrite Eq. (64) in terms of $a_{L},\,b_{L}$ and $c_{L}$,
$|P_{\rm tr}(x,t)\rangle$$\approx$ $\frac{1}{2}\bigg{(}2\,a_{L}(x_{0},x)\begin{bmatrix}1\\ 1\end{bmatrix}+~{}m_{0}\,\frac{v}{\omega}\,b_{L}(x,x_{0})\begin{bmatrix}1\\ 1\end{bmatrix}-\frac{v}{\omega}\,b_{L}(x_{0},x)\begin{bmatrix}1\\ -1\end{bmatrix}$ (B.7)
@@
$+~{}\frac{v}{\omega}\,\big{(}1-\frac{v}{\sqrt{2\omega D_{e}}}\big{)}\bigg{\{}e^{-k_{a}x}\,c_{L}(x_{0})\begin{bmatrix}1\\ a^{(1)}\end{bmatrix}+e^{-k_{a}(L-x)}\,c_{L}(L-x_{0})\begin{bmatrix}a^{(1)}\\ 1\end{bmatrix}\bigg{\}}\bigg{)},$
which gives an approximate closed form expression of the propagator in the large $L$ limit in terms of known functions. We obtain Eq. (65) in the limit $L\rightarrow\infty$. It is evident that unlike in each of the eigenstates, the ‘active’ contributions appear at the leading order in the general time-dependent distribution.
Appendix C Large time distribution of a passive underdamped Brownian motion in presence of an absorbing barrier
@@
$\widetilde{P}_{x_{0},v_{0}}(x,v;s)$ $=$ $\frac{e^{-v^{2}/2}}{\sqrt{8\pi}}\,\sum_{n=0}^{\infty}\frac{e^{-q_{n}|x-x_{0}|}}{q_{n}}\,f_{n}^{\mp}(v)f_{n}^{\mp}(v_{0})$ (C.1)
@@
where $q_{n}=\sqrt{n+s},~{}\sigma_{m\,n}=\frac{1}{q_{m}+q_{n}}\,\frac{1}{Q_{m}\,Q_{n}}=\sigma_{n\,m}$ with $Q_{n}=\lim_{N\rightarrow\infty}\frac{\sqrt{n!N!}\,\exp(2q_{n}\sqrt{N+1})}{\prod_{r=0}^{N+n}(q_{r}+q_{n})}$, and $f_{n}^{+}(v)=f_{n}(v)=\frac{e^{v^{2}/4}}{\sqrt{n!}}\left[(-1)^{n}e^{z^{2}/4}\frac{d^{n}}{dz^{n}}e^{-z^{2}/2}\right]_{z=2q_{n}-v}\,$, $f_{n}^{-}(v)=f_{n}(-v)$. Here all the variables are non-dimensionalised, $t\rightarrow t\,\gamma/m,~{}x\rightarrow x\,\sqrt{\gamma/(mD)}\,,~{}v\rightarrow v\,\sqrt{m/(D\gamma)}\,$, $m$ is the mass of the particle. To find the distribution at large times we need to extract the leading behaviour of $\widetilde{P}$ in the $s\rightarrow 0$ limit. In this limit, the summations with only $m=0$ or $n=0$ will contribute in Eq. (C.1), and the different quantities are evaluated as,
@@
[…]
@@
$\widetilde{P}_{x_{0},v_{0}}(x;s)\approx\frac{e^{-\frac{3}{2}s}}{2\,\sqrt{s}}\,\left[e^{-(|x-x_{0}|\pm v_{0})\sqrt{s}}-e^{-(x+x_{0}+2\,l_{M}+v_{0})\sqrt{s}}\right]+\widetilde{P}_{BL}(x;s),$ (C.2)
where $l_{M}\approx 1.458$, ±$\pm~{}{\rm is~{}for~{}}x\lessgtr x_{0}$, and the boundary layer profile is given by,
$\widetilde{P}_{BL}(x;s)$$\approx$ $\frac{1}{2}\sum_{n=0}^{\infty}\left[\frac{1+\sqrt{n}}{n}\right]^{\frac{1}{2}}\frac{e^{-\frac{3}{2}n-\sqrt{n}}}{n!\,b_{n}}e^{-(v_{0}-\frac{1}{\sqrt{n}})\sqrt{s}}$ (C.3)
×$\times\left[H(n)\,e^{-x\sqrt{n}}\,e^{-(x_{0}+v_{0})\sqrt{s}}+\Phi_{n}(-v_{0})\,e^{-(x_{0}-v_{0})\sqrt{n}}\,e^{-x\sqrt{s}}\right],$
[…]
The distribution is markedly different from that obtained in the usual overdamped problem. It contains a rich ‘kinetic boundary layer’ structure of a finite skin depth $\sim l_{p}=\sqrt{mD/\gamma}$ near the absorbing walls. Secondly, far from the wall at large times the distribution takes a scaling-like form $P_{\rm sc}=P_{x_{0},v_{0}}^{\infty}(x,t)-P_{\rm BL}$ which corresponds to a spatial shift $x\rightarrow x+l_{M}$. $P_{\rm sc}$ satisfies the absorbing condition at $x=-l_{M}$ instead of $x=0$ [54], i.e. $P_{\rm sc}(-l_{M},t)=0$, implying that $l_{M}$ is the Milne length. In [64] the Milne length is exactly determined, $l_{M}=-\zeta(\frac{1}{2})\,l_{p}\approx 1.460\,l_{p}\,$, a value approximated remarkably well by the Euler-Maclaurin formula. Also note the factor $e^{-\frac{3}{2}s}$ in Eq. (C.2) that corresponds to a temporal shift, $t\rightarrow t-\frac{3}{2}\tau_{p}$ with $\tau_{p}=m/\gamma$, suggesting that $P_{\rm sc}$ pertains to an initial condition at $t=\frac{3}{2}\tau_{p}$ instead of $t=0$ [54].
Appendix D Distribution of ABPs on a semi-infinite line with absorbing condition at $x=0$

### v4

@@ Appendix B Exit probability of RTP at the left boundary
$E_{m}(x)=\frac{(1+e^{-\mu L})\,x-\frac{v}{2\,\omega}(m-\frac{v}{\mu\,D})\,[1-e^{-\mu x}]+\frac{v}{2\,\omega}(m+\frac{v}{\mu\,D})\,[e^{-\mu(L-x)}-e^{-\mu L}]}{L\,(1+e^{-\mu L})+\frac{v^{2}}{\omega\,\mu\,D}\,(1-e^{-\mu L})}.$ (B.2)
@@
Appendix C Subleading correction to the band near $\lambda=0$
We want to solve Eq. (47) to find the leading $L$-dependence of $\lambda_{n}^{(1)}$. Putting $\lambda_{n}=\lambda_{n}^{(0)}+\lambda_{n}^{(1)}$ in Eq. (47), we get,
$e^{-L\sqrt{\frac{\lambda_{n}^{(0)}+\lambda_{n}^{(1)}}{D_{e}}}}\approx\mp\left(1+\frac{v^{2}}{\omega D_{e}}\sqrt{\frac{\lambda_{n}^{(0)}+\lambda_{n}^{(1)}}{2\omega}}\right).$ (C.1)
To the first subleading order the left hand side of the Eq. (C.1) can be written as,
$e^{-L\sqrt{\frac{\lambda_{n}^{(0)}+\lambda_{n}^{(1)}}{D_{e}}}}$ $=$ $e^{-L\sqrt{\frac{\lambda_{n}^{(0)}}{D_{e}}}\,\left(1+\frac{\lambda_{n}^{(1)}}{\lambda_{n}^{(0)}}\right)^{1/2}}\approx e^{-L\sqrt{\frac{\lambda_{n}^{(0)}}{D_{e}}}}~e^{-L\sqrt{\frac{\lambda_{n}^{(0)}}{D_{e}}}\,\left(\frac{\lambda_{n}^{(1)}}{2\,\lambda_{n}^{(0)}}\right)}$ (C.2)
Now, using in the above the expression for $\lambda_{n}^{(0)}$ from Eq. (45), the l.h.s. of Eq. (C.1) is obtained as,
$e^{-L\sqrt{\frac{\lambda_{n}^{(0)}+\lambda_{n}^{(1)}}{D_{e}}}}\approx\mp 1\times\exp\left(-i\,\frac{n\pi}{2}\,\frac{\lambda_{n}^{(1)}}{\lambda_{n}^{(0)}}\right)\approx\mp 1\pm i\,\frac{n\pi}{2}\,\frac{\lambda_{n}^{(1)}}{\lambda_{n}^{(0)}},$ (C.3)
Similarly, we find the right hand side of Eq. (C.1) to the first subleading order,
Then, equating the l.h.s. and r.h.s. of Eq. (C.1), we get Eq. (48) as the leading correction.
Appendix D Calculation of the coefficients $\beta^{(i)}$, expression for the left eigenvectors of $\mathcal{L}$ in the $\lambda=0$ band and the distribution for large but finite $L$
In section 4.4 we write the form of the distribution for large but finite systems with size $L$. Here we first evaluate the coefficients $\beta^{(i)}$. Eqs. (52)-(54) gives the relation among the coefficients,
$\beta^{(1)}$ $=$ $-\,\beta^{(3)}\,\frac{a^{(3)}\pm e^{k_{b}\,L}}{a^{(1)}\pm e^{k_{a}\,L}},$ (D.1)
$\beta^{(2)}$ $=$ $\pm\,\beta^{(1)}a^{(1)}\,e^{k_{a}\,L},$ (D.2)
$\beta^{(4)}$ $=$ $\pm\,\beta^{(3)}a^{(3)}\,e^{k_{b}\,L},$ (D.3)
In the large $L$ limit we find, for $o_{x}=\pm 1$, $e^{k_{b}\,L}\approx\mp(1-\frac{n\,\pi}{L}\,\frac{v^{2}}{\omega\,\sqrt{2\omega D_{e}}})$ using Eqs. (C.3) and (48), and $\beta^{(1)}=-\,\beta^{(3)}\,\frac{a^{(3)}\pm\,e^{k_{b}\,L}}{a^{(1)}\pm\,e^{k_{a}\,L}}\approx\mp\,\beta^{(3)}\,[a^{(3)}\pm e^{k_{b}\,L}]\,e^{-k_{a}\,L}$ from Eq. (D.1). In Eq. (27), keeping the first subleading term in the expression for $a^{(3)}$ and using Eq. (49) to determine $k_{b}=\sqrt{\lambda/D_{e}}$ to $O(L^{-1})$, we find, $a^{(3)}\approx 1-vk_{b}/\omega\approx 1-i\,\frac{v}{\omega}\,\frac{n\pi}{L}$. Putting these together in Eq. (D.1) we obtain,
$\beta^{(1)}=\pm\,i\,\beta^{(3)}\,\frac{n\pi}{L}\,\frac{v}{\omega}\big(1-\frac{v}{\sqrt{2\omega D_{e}}}\big)\,e^{-k_{a}\,L}\,,{~\rm for~}n={\rm odd/even}.$ (D.4)
Using the above in Eq. (D.2) we find,
$\beta^{(2)}=\,i\,\beta^{(3)}\,\frac{n\pi}{L}\,a^{(1)}\,\frac{v}{\omega}\big(1-\frac{v}{\sqrt{2\omega D_{e}}}\big);$ (D.5)
and finally, from Eq. (D.3)
$\beta^{(4)}=-\beta^{(3)}a^{(3)}\big(1-i\,\frac{n\pi}{L}\,\frac{v^{2}}{\omega\sqrt{2\omega D_{e}}}\big).$ (D.6)
Using the coefficients in Eq. (28), the eigenvector $|\phi_{n}(x)\rangle$ is found up to a prefactor $\beta_{n}^{(3)}$ (we resume the suffix $n$). To determine $\beta_{n}^{(3)}$ we use an orthonormality condition, for which we need to find the left eigenvectors $\langle\psi_{n}|$ of $\mathcal{L}$. Noting that it is just the transpose of the right eigenvector of $\mathcal{L}^{\dagger}\equiv\mathcal{L}|_{\partial_{x}\rightarrow-\partial_{x}}$ we write:
$\langle\psi_{n}(x)|$ $=$ $|\phi_{n}(x,-v)\rangle^{T}\approx\tilde{\beta}_{n}^{(3)}\bigg(2\,i\,\sin\frac{n\pi x}{L}\begin{bmatrix}1\\ 1\end{bmatrix}^{T}+\,i\,\frac{n\pi v}{\omega\,L}\cos\frac{n\pi x}{L}\begin{bmatrix}1\\ -1\end{bmatrix}^{T}$ (D.7)
@@ Appendix C Subleading correction to the band near $\lambda=0$
$\Rightarrow\int_{0}^{L}dx{\langle\psi_{n}(x)|\phi_{m}(x)\rangle}=-4\,\tilde{\beta}^{(3)}_{n}\beta^{(3)}_{n}L\,{\left(1-i(i+n\pi)\,\frac{2\,l_{M}}{L}\right)}\,\delta_{mn}~,~l_{M}=\frac{v^{2}}{2\omega\,\sqrt{2\omega D_{e}}}$ (D.8)
@@
$\tilde{\beta}^{(3)}_{n}\beta_{n}^{(3)}\approx-\frac{1}{4L}\left(1+i(i+n\pi)\,\frac{2\,l_{M}}{L}\right).$ (D.9)
The knowledge of the $\tilde{\beta}\beta$ product is sufficient to determine the large time distribution, which we find next. For the initial condition, $|P(x,0)\rangle=\delta(x-x_{0})\,\begin{bmatrix}s_{+}\\ s_{-}\end{bmatrix}$, we find,
$\begin{split}c_{n}&=\int_{0}^{L}dx\,\langle\psi_{n}(x)|P(x,0)\rangle\approx\tilde{\beta}^{(3)}_{n}\bigg[2i\sin\frac{n\pi x_{0}}{L}+\frac{in\pi v}{\omega L}(s_{+}-s_{-})\cos\frac{n\pi x_{0}}{L}\\ &~~~~~-\frac{n\pi v}{\omega L}(1-\frac{v}{\sqrt{2\omega D_{e}}})\sin\frac{n\pi x_{0}}{L}+\frac{in\pi v^{2}}{\omega L\sqrt{2\omega D_{e}}}(1-\frac{2x_{0}}{L})\cos\frac{n\pi x_{0}}{L}\\ &~~~~~-\frac{in\pi v}{\omega L}(1+\frac{v}{\sqrt{2\omega D_{e}}})\bigg\{(-1)^{n-1}e^{-k_{a}(L-x_{0})}(\tilde{a}_{n}^{(1)}s_{+}+s_{-})+e^{-k_{a}x_{0}}(s_{+}+\tilde{a}_{n}^{(1)}s_{-})\bigg\}\bigg].\end{split}$ (D.10)
Using Eqs. (D.9) and (D.10) in Eq. (56), we obtain Eq. (58) for the propagator to $O(L^{-1})$ for times $\omega^{-1}\ll t\ll L^{2}/D_{e}$. To find a closed-form expression, we define the following quantities corresponding to the summations that occur in Eq. (58) and keep only the leading terms after using the Poisson summation formula:
@@
The $L$-dependent terms in the square brackets, which are very small for $L\gg x_{1},x_{2}$, are kept in view of the symmetries satisfied by the discrete summation. These also preserve the symmetries of the density and magnetisation profiles:
$\begin{split}P(x,t;x_{0},m_{0})&=P(L-x,t;L-x_{0},-m_{0}),\\ Q(x,t;x_{0},m_{0})&=-Q(L-x,t;L-x_{0},-m_{0}).\end{split}$ (D.12)
Using Eq. (D.11) we rewrite Eq. (58) in terms of $a_{L},\,b_{L}$ and $c_{L}$,
$|P_{\rm tr}(x,t)\rangle$ $\approx$ $\frac{1}{2}\bigg(2\,a_{L}(x_{0},x)\begin{bmatrix}1\\ 1\end{bmatrix}+~m_{0}\,\frac{v}{\omega}\,b_{L}(x,x_{0})\begin{bmatrix}1\\ 1\end{bmatrix}-\frac{v}{\omega}\,b_{L}(x_{0},x)\begin{bmatrix}1\\ -1\end{bmatrix}$ (D.13)
@@
[…]
which gives an approximate closed form expression of the propagator in the large $L$ limit in terms of known functions. We obtain Eq. (59) in the limit $L\rightarrow\infty$. It is evident that unlike individual eigenstates, the ‘active’ contributions appear at the leading order in the time-dependent distribution.
Appendix E Large time distribution of a passive underdamped Brownian motion in presence of an absorbing barrier
@@
$\widetilde{P}_{x_{0},v_{0}}(x,v;s)$ $=$ $\frac{e^{-v^{2}/2}}{\sqrt{8\pi}}\,\sum_{n=0}^{\infty}\frac{e^{-q_{n}|x-x_{0}|}}{q_{n}}\,f_{n}^{\mp}(v)f_{n}^{\mp}(v_{0})$ (E.1)
@@
where $x_{0},v_{0}$ are the initial positions and velocities respectively, the absorbing boundary is at $x=0$, $q_{n}=\sqrt{n+s},~\sigma_{m\,n}=\frac{1}{q_{m}+q_{n}}\,\frac{1}{Q_{m}\,Q_{n}}=\sigma_{n\,m}$ with $Q_{n}=\lim_{N\rightarrow\infty}\frac{\sqrt{n!N!}\,\exp(2q_{n}\sqrt{N+1})}{\prod_{r=0}^{N+n}(q_{r}+q_{n})}$, and $f_{n}^{+}(v)=f_{n}(v)=\frac{e^{v^{2}/4}}{\sqrt{n!}}\left[(-1)^{n}e^{z^{2}/4}\frac{d^{n}}{dz^{n}}e^{-z^{2}/2}\right]_{z=2q_{n}-v}\,$, $f_{n}^{-}(v)=f_{n}(-v)$. Here all the variables are nondimensionalised, $t\rightarrow t\,\gamma/m,~x\rightarrow x\,\sqrt{\gamma/(mD)}\,,~v\rightarrow v\,\sqrt{m/(D\gamma)}\,$, $m$ is the mass of the particle, $\gamma$ is the damping coefficient, and $D$ is the thermal diffusivity. To find the distribution at large times we need to extract the leading behaviour of $\widetilde{P}$ in Eq. (E.1) in the $s\rightarrow 0$ limit. In this limit, the terms in the summations with only $m=0$ or $n=0$ will contribute, and the different quantities are evaluated as,
@@
    The $Q_{n}$’s are evaluated using Euler-Maclaurin formula: $\sum_{r=1}^{N}f(r)=\int_{1}^{N}f(x)dx+\frac{f(1)+f(N)}{2}+\mathcal{R}_{N}$, where $\mathcal{R}_{N}$ is the residue. Without considering $\mathcal{R}_{N}$ we find,
@@
$\widetilde{P}_{x_{0},v_{0}}(x;s)\approx\frac{e^{-\frac{3}{2}s}}{2\,\sqrt{s}}\,\left[e^{-(|x-x_{0}|\pm v_{0})\sqrt{s}}-e^{-(x+x_{0}+2\,l_{M}+v_{0})\sqrt{s}}\right]+\widetilde{P}_{BL}(x;s),$ (E.2)
where $l_{M}\approx 1.458$, $\pm~{\rm is~for~}x\lessgtr x_{0}$, and
$\widetilde{P}_{BL}(x;s)$ $\approx$ $\frac{1}{2}\sum_{n=0}^{\infty}\left[\frac{1+\sqrt{n}}{n}\right]^{\frac{1}{2}}\frac{e^{-\frac{3}{2}n-\sqrt{n}}}{n!\,b_{n}}e^{-(v_{0}-\frac{1}{\sqrt{n}})\sqrt{s}}$ (E.3)
$\times\left[H(n)\,e^{-x\sqrt{n}}\,e^{-(x_{0}+v_{0})\sqrt{s}}+\Phi_{n}(-v_{0})\,e^{-(x_{0}-v_{0})\sqrt{n}}\,e^{-x\sqrt{s}}\right],$
with $H(n)=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty}dv\,\Phi_{n}(v)\,e^{-\frac{v^{2}}{2}+v\,\sqrt{n}}$ . Plugging in the units and taking the inverse transform, we find the leading time dependence of the density profile for the problem,
${P}_{x_{0},v_{0}}(x;t)$ $\approx$ $\frac{1}{\sqrt{4\pi D_{e}t}}\left[e^{-\frac{(|x-x_{0}|\,\pm\,v_{0}\,m/\gamma)^{2}}{4Dt}}-e^{-\frac{(x+x_{0}+2\,l_{M}\,+\,v_{0}\,m/\gamma)^{2}}{4Dt}}\right]+P_{BL}(x,t),$ (E.4)
where the boundary layer term $P_{BL}(x,t)$ is given by the inverse Laplace transform of Eq. (E.3). Quite remarkably, for large $t$, expanding the Gaussian terms in Eq. (E.4) around $|x-x_{0}|$ and $(x+x_{0})$ respectively and keeping the terms up to $O(xt^{-3/2})$ reduces the density to a form similar to Eq. (60) obtained for RTP, except the boundary layer.
Note that the distribution is markedly different from that obtained in the overdamped problem. It contains a rich ‘kinetic boundary layer’ structure of a finite skin depth $\sim l_{p}=\sqrt{mD/\gamma}$ near the absorbing walls. Secondly, far from the wall at large times the distribution takes a scaling-like form $P_{\rm sc}=P_{x_{0},v_{0}}^{\infty}(x,t)-P_{\rm BL}$ which corresponds to a spatial shift $x\rightarrow x+l_{M}$ and $x_{0}\rightarrow x_{0}+l_{M}+v_{0}\,m/\gamma$. $P_{\rm sc}$ satisfies the absorbing condition at $x=-l_{M}$ instead of $x=0$ [55], i.e. $P_{\rm sc}(-l_{M},t)=0$, implying that $l_{M}$ is the Milne length. In [67] the Milne length is exactly determined, $l_{M}=-\zeta(\frac{1}{2})\,l_{p}\approx 1.460\,l_{p}\,$, a value approximated remarkably well by the Euler-Maclaurin formula. Also note the factor $e^{-\frac{3}{2}s}$ in Eq. (E.2) that corresponds to a temporal shift, $t\rightarrow t-\frac{3}{2}\tau_{p}$ with $\tau_{p}=m/\gamma$, suggesting that $P_{\rm sc}$ pertains to an initial condition at $t=\frac{3}{2}\tau_{p}$ instead of $t=0$ [55].
Appendix F Distribution of ABPs on a semi-infinite line with absorbing condition at $x=0$